# Learning Equilibria from Data: Provably Efficient Multi-Agent Imitation Learning

### v1

@@ Appendix I Extension to $n$-player general-sum Games
K.1 Environments
We consider two different environments for our numerical validation, one that has ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})<\infty,$ and the lower bound construction Fig. 1 with different NE experts to control ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}}).$ In particular we have multiple with ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})<\infty$ and the same NE expert as in Theorem 3.2 to get ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=\infty.$
@@
For this we consider two environments. For the first environment, we generate a random Zero-Sum Markov Game with $\left|{{\mathcal{S}}}\right|=10,\left|{{\mathcal{A}}}\right|=\left|{{\mathcal{B}}}\right|=3$ and a reward between $-1$ and $1$. To ensure that the expert covers all states we use a uniform initial state distribution, i.e $d_{0}(s_{0}):=\mathrm{Unif}({\mathcal{S}}).$ We set the discount factor to $0.9$
Additionally, we choose the Markov Game from the Lower bound construction and use that the set of Nash equilibria is convex for Zero-sum Games. This way we take a mixture of Nash equilibria that chooses the $S_{\mathrm{copy}}$ path and the blue path, for a detailed description see Section K.2.
@@
For ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=\infty$, we use the Zero-Sum Markov Game given in Fig. 1 with the simplification that $\left|{S_{\mathrm{xplt1}}}\right|=\left|{S_{\mathrm{xplt1}}}\right|=\left|{S_{\mathrm{copy}}}\right|=1$ as our goal here is only to verify the theoretical insights, but not to prove that also the transition model cannot be used for non-interactive Imitation Learning algorithms. This means that we have $\left|{{\mathcal{S}}}\right|=7.$ Additionally, we have $\left|{{\mathcal{A}}}\right|=\left|{{\mathcal{B}}}\right|=3$ and $d_{0}(s_{0})=\delta_{s_{0}}.$ We set the discount factor to $0.9$ The reward is given by $R(S_{\mathrm{xplt2}}),$ $R(S_{\mathrm{xplt1}})=-0.1$ and 00 otherwise.
@@
To get control over $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$, note that the set of Nash equilibria is convex for Two Player Zero-Sum Games. Therefore, consider the Lower bound example illustrated in Fig. 1. Note that we have a pure NE that chooses action $a_{3}b_{3}$ to get on the blue path. Choosing a different pure action, let us assume $a_{2},b_{2}$ will lead the agent to choose the path that goes on $s_{1},S_{\mathrm{copy}}.$ Now, as the set of NE is convex, we can also mix these equilibria. To choose the minimal $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}}),$ we pick for $(a)$ the Nash equilibrium such that $\mu^{\operatorname{E}}(a_{2}\mid s_{0})=\nu^{\operatorname{E}}(a_{2}\mid s_{0})=\mu^{\operatorname{E}}(a_{3}\mid s_{0})=\nu^{\operatorname{E}}(a_{3}\mid s_{0})=0.5.$ To increase $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}}),$ we have to increase the probability of the experts to take action $a_{3}$ and $b_{3}$ respectively. We choose for $(c)$ $\mu^{\operatorname{E}}(a_{3}\mid s_{0})=\nu^{\operatorname{E}}(a_{3}\mid s_{0})=0.999,$ for $(c)$ we pick $\mu^{\operatorname{E}}(a_{3}\mid s_{0})=\nu^{\operatorname{E}}(a_{3}\mid s_{0})=0.9999.$ For $(d),$ we use the same expert policy as in the lower bound construction, i.e. $\mu^{\operatorname{E}}(a_{3}\mid s_{0})=\nu^{\operatorname{E}}(a_{3}\mid s_{0})=1.$
@@
K.3 Practical considerations
Next, we list practical considerations for our algorithms, that could speed up the performance. First, note that while solving the RL inner loop in Algorithm 2 can be computationally expensive, the objective between successive iterations changes only through the updates of the policies $\mu_{k}$ and $\nu_{k}$. Consequently, if these policies change only slightly between iterations, the optimal solutions for $y_{k}$ and $z_{k}$ may also vary only marginally. This observation suggests that initializing the optimization with the solution from the previous iteration, a common technique known as warm-start optimization, can significantly accelerate convergence.
@@
Last, note that we assumed for our analysis that there is no initial dataset for interactive Imitation Learning Section 2. However, in general it is possible to consider an initial dataset ${\mathcal{D}}$, from which we can learn initial policies with a non-interactive Imitation Learning algorithm like BC. This can speed up the convergence of our proposed algorithms as the maximum uncertainty exploration will mainly focus on states out of the distribution from the initial dataset. We give the algorithm of MURMAIL with an initial dataset in Algorithm 4. Similarly, one can adjust Algorithm 1.
@@
K.4 Additional plots
In this section, we list an additional plot for the second more involved environment that has $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})<\infty.$ Here we can observe similarly to case $(a)$ of Fig. 2 that the speed of convergence from BC is higher compared to MURMAIL. It indicates that the chosen algorithm has a small concentrability coefficient ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ and again highlights the importance of algorithm selection depending on the underlying environment.
@@
Appendix L Useful Results
@@
Lemma L.1 (see e.g. Lemma IX.5 by Alatur et al. [2024]).
@@
Lemma L.2 (Concentration Inequality for Total Variation Distance, see e.g. Thm 2.1 by Berend and Kontorovich [2012]).
@@
By McDiarmid’s inequality, we have for any $\epsilon>0$,
$\Pr\left(f-\mathbb{E}[f]\geq\epsilon\right)\leq\exp\left(-\frac{n\epsilon^{2}}{2}\right).$
@@
$\exp\left(-\frac{n\epsilon^{2}}{2}\right)=\delta\quad\implies\quad\epsilon=\sqrt{\frac{2\log(1/\delta)}{n}}.$
@@
Lemma L.3 (Binomial concentration, see e.g. Lemma A.1 by Xie et al. [2021]).
@@
We consider two cases. Case 1: $p\leq\frac{8\log(1/\delta)}{n}.$ As $N\vee 1\geq 1,$ we have $\frac{p}{N\vee 1}\leq p\leq\frac{8\log(1/\delta)}{n}$ almost surely. Case 2: $p>\frac{8\log(1/\delta)}{n}.$ Note, that then $\mathbb{E}[N]=np>8\log(1/\delta)$ and by the multiplicative Chernoff bound, for any $10<\epsilon<$ 1 it holds true that
$\mathbb{P}\left(N<(1-\epsilon)np\right)\leq\exp\left(-\frac{\epsilon^{2}}{2}np\right).$
Now, with $\epsilon=\frac{1}{2}$ we have
$\mathbb{P}\left(N<(1-\epsilon)np\right)\leq\exp\left(-\frac{np}{8}\right)\leq\delta.$

### v2

@@ Appendix H Extension to $n$-player general-sum games
J.1 Environments
We consider two different environments for our numerical validation, one that has ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})<\infty,$ and the lower bound construction Fig.˜1 with different NE experts to control ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}}).$ In particular we have multiple with ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})<\infty$ and the same NE expert as in Theorem˜3.2 to get ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=\infty.$
@@
For this we consider two environments. For the first environment, we generate a random Zero-Sum Markov Game with $\left|{{\mathcal{S}}}\right|=10,\left|{{\mathcal{A}}}\right|=\left|{{\mathcal{B}}}\right|=3$ and a reward between 1-1 and $1$. To ensure that the expert covers all states we use a uniform initial state distribution, i.e $d_{0}(s_{0}):=\mathrm{Unif}({\mathcal{S}}).$ We set the discount factor to $0.9.$
Additionally, we choose the Markov Game from the Lower bound construction and use that the set of Nash equilibria is convex for Zero-sum Games. This way we take a mixture of Nash equilibria that chooses the $S_{\mathrm{copy}}$ path and the blue path, for a detailed description see Section˜J.2.
@@
For ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=\infty$, we use the Zero-Sum Markov Game given in Fig.˜1 with the simplification that $\left|{S_{\mathrm{xplt1}}}\right|=\left|{S_{\mathrm{xplt1}}}\right|=\left|{S_{\mathrm{copy}}}\right|=1$ as our goal here is only to verify the theoretical insights, but not to prove that also the transition model cannot be used for non-interactive Imitation Learning algorithms. This means that we have $\left|{{\mathcal{S}}}\right|=7.$ Additionally, we have $\left|{{\mathcal{A}}}\right|=\left|{{\mathcal{B}}}\right|=3$ and $d_{0}(s_{0})=\delta_{s_{0}}.$ We set the discount factor to $0.9.$ The reward is given by $R(S_{\mathrm{xplt2}}),$ $R(S_{\mathrm{xplt1}})=-0.1$ and 0 otherwise.
@@
To get control over $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$, note that the set of Nash equilibria is convex for Two Player Zero-Sum Games. Therefore, consider the Lower bound example illustrated in Fig.˜1. Note that we have a pure NE that chooses action $a_{3}b_{3}$ to get on the blue path. Choosing a different pure action, let us assume $a_{2},b_{2}$ will lead the agent to choose the path that goes on $s_{1},S_{\mathrm{copy}}.$ Now, as the set of NE is convex, we can also mix these equilibria. To choose the minimal $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}}),$ we pick for $(a)$ the Nash equilibrium such that $\mu^{\operatorname{E}}(a_{2}\mid s_{0})=\nu^{\operatorname{E}}(a_{2}\mid s_{0})=\mu^{\operatorname{E}}(a_{3}\mid s_{0})=\nu^{\operatorname{E}}(a_{3}\mid s_{0})=0.5.$ To increase $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}}),$ we have to increase the probability of the experts to take action $a_{3}$ and $b_{3}$ respectively. We choose for $(c)$ $\mu^{\operatorname{E}}(a_{3}\mid s_{0})=\nu^{\operatorname{E}}(a_{3}\mid s_{0})=0.999,$ for $(c)$ we pick $\mu^{\operatorname{E}}(a_{3}\mid s_{0})=\nu^{\operatorname{E}}(a_{3}\mid s_{0})=0.9999.$ For $(d),$ we use the same expert policy as in the lower bound construction, i.e. $\mu^{\operatorname{E}}(a_{3}\mid s_{0})=\nu^{\operatorname{E}}(a_{3}\mid s_{0})=1.$
@@
J.3 Practical considerations
Next, we list practical considerations for our algorithms, that could speed up the performance. First, note that while solving the RL inner loop in Algorithm˜2 can be computationally expensive, the objective between successive iterations changes only through the updates of the policies $\mu_{k}$ and $\nu_{k}$. Consequently, if these policies change only slightly between iterations, the optimal solutions for $y_{k}$ and $z_{k}$ may also vary only marginally. This observation suggests that initializing the optimization with the solution from the previous iteration, a common technique known as warm-start optimization, can significantly accelerate convergence.
@@
Last, note that we assumed for our analysis that there is no initial dataset for interactive Imitation Learning Section˜2. However, in general it is possible to consider an initial dataset $\mathcal{D}$}, from which we can learn initial policies with a non-interactive Imitation Learning algorithm like BC. This can speed up the convergence of our proposed algorithms as the maximum uncertainty exploration will mainly focus on states out of the distribution from the initial dataset. We give the algorithm of MURMAIL with an initial dataset in Algorithm˜4. Similarly, one can adjust Algorithm˜1.
@@
J.4 Additional plots
In this section, we list an additional plot for the second more involved environment that has $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})<\infty.$ Here we can observe similarly to case $(a)$ of Fig.˜2 that the speed of convergence from BC is higher compared to MURMAIL. It indicates that the chosen algorithm has a small concentrability coefficient ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ and again highlights the importance of algorithm selection depending on the underlying environment.
@@
Appendix K Useful Results
@@
Lemma K.1 (see e.g. Lemma IX.5 by Alatur et al. [2024]).
@@
Lemma K.2 (Concentration Inequality for Total Variation Distance, see e.g. Thm 2.1 by Berend and Kontorovich [2012]).
@@
By McDiarmid’s inequality, we have for any $\varepsilon>0$,
$\Pr\left(f-\mathbb{E}[f]\geq\varepsilon\right)\leq\exp\left(-\frac{n\varepsilon^{2}}{2}\right).$
@@
$\exp\left(-\frac{n\varepsilon^{2}}{2}\right)=\delta\quad\implies\quad\varepsilon=\sqrt{\frac{2\log(1/\delta)}{n}}.$
@@
Lemma K.3 (Binomial concentration, see e.g. Lemma A.1 by Xie et al. [2021]).
@@
We consider two cases. Case 1: $p\leq\frac{8\log(1/\delta)}{n}.$ As $N\vee 1\geq 1,$ we have $\frac{p}{N\vee 1}\leq p\leq\frac{8\log(1/\delta)}{n}$ almost surely. Case 2: $p>\frac{8\log(1/\delta)}{n}.$ Note, that then $\mathbb{E}[N]=np>8\log(1/\delta)$ and by the multiplicative Chernoff bound, for any $0<\varepsilon<1$ it holds true that
$\mathbb{P}\left(N<(1-\varepsilon)np\right)\leq\exp\left(-\frac{\varepsilon^{2}}{2}np\right).$
Now, with $\varepsilon=\frac{1}{2}$ we have
$\mathbb{P}\left(N<(1-\varepsilon)np\right)\leq\exp\left(-\frac{np}{8}\right)\leq\delta.$